OpenStudy (anonymous):

What is the next number in the following pattern? 1, 3, 7, 15, 31, , … A. 65 B. 33 C. 63 D. 47

OpenStudy (trojanpoem):

63

OpenStudy (anonymous):

ok but can you explain

OpenStudy (amoodarya):

\[1\\3=1+2^1\\7=3+2^2\\15=7+2^3\\31=15+2^4\\?\\31+2^5=31+32=63\]

OpenStudy (trojanpoem):

3 - 1= 2 7 -3 = 4 15 - 7 = 8 31 - 15 = 16 so we multiply by 2 ( 16/ 8 = 2) , 4/2 = 2 so 16 * 2 = 32 + 31 = 63

OpenStudy (anonymous):

tro gve amood a medal

OpenStudy (anonymous):

give and thx yall

OpenStudy (anonymous):

wait one more

OpenStudy (anonymous):

2, -3, 4, -5, _____ A. 3 B. 6 C. -5 D. 4 E. -6

OpenStudy (amoodarya):

\[a_{n+1}=2a_n+1\\a_1=1\]

OpenStudy (anonymous):

so its -6

OpenStudy (amoodarya):

Tower of Hanoi

OpenStudy (anonymous):

/???

OpenStudy (amoodarya):

it is the name of first question for the second 2,3,4,5,6,7,8,... \[a_n=n+1\] 2,-3,4,-5,+6,-7,+8 \[a_n=(-1)^{n+1}(n+1)\]

OpenStudy (anonymous):

itts -6

OpenStudy (anonymous):

i mean 6

OpenStudy (anonymous):

lol